OpenStudy (anonymous):

in 2D vector the perpendicular of ( 2 3) is - ( 3 2) What about in 3D vector?

OpenStudy (anonymous):

\[\left(\begin{matrix}2 \\ 3\end{matrix}\right) become -\left(\begin{matrix}3 \\ 2\end{matrix}\right)\] what about in 3D vector?

Parth (parthkohli):

There are many perpendicular vectors.

OpenStudy (anonymous):

so those inverses only apply at 2D vector? (as it contain only on perpendicular)

OpenStudy (zzr0ck3r):

(2,3) is not a vector in 3 space

OpenStudy (anonymous):

ok thanks